# On the Diameter of Arrangements of Topological Disks

### v1

@@ 3 The general case of $n$ disks
Recall that $\mu(A)$ denotes the number of maximal faces in the arrangement $\mathcal{A}$ defined by the set $\mathcal{D}$ of $n$ disks. To bound $\mu(\mathcal{A})$, we will bound the number of maximal faces that are contained in a fixed disk from $\mathcal{D}$, say $D_{0}$, and that are incident to $\partial D_{0}$. Multiplying this number by $n$ then gives an upper bound on $\mu(A)$.
The intersection points $\partial D_{0}\cap\partial D_{i}$, with $i\neq 0$, partition $\partial D_{0}$ into a number of elementary intervals, [^1] We ignore the trivial case where $\partial D_{0}$ is not intersected by any of the other boundaries $\partial D_{i}$. as shown in Fig. 5.
Figure 5: The partitioning of $\partial D_{0}$ into elementary intervals (indicated in black). The colored intervals (which are drawn slightly outside $\partial D_{0}$ for clarity) indicate the intersections of $\partial D_{0}$ with the various components in $\mathcal{C}(D_{0},D_{i})$, where intervals of the same color belong to the same component. The different shades of green indicate that the intervals belong to different components of the green disk and the blue disk $D_{0}$.
We denote the set of these elementary intervals by $\mathcal{I}$. Recall that $\mathcal{C}(D_{i},D_{j})$ denotes the set of connected components of $D_{i}\cap D_{j}$. Define $\mathcal{C}(\mathcal{D})\coloneqq\bigcup_{0\leqslant i<j<n}\mathcal{C}(D_{i},D_{j})$ to be the set of all such connected components, which we will simply call components from now on. For an elementary interval $I\in\mathcal{I}$, define
$\Lambda(I)\coloneqq\{C\in\mathcal{C}(\mathcal{D}):I\subseteq C\}$
to be the component set of $I$. Furthermore, we define $f(I)$ to be the face of $\mathcal{A}(\mathcal{D})$ that is incident to $I$ and lies inside $D_{0}$. We call $f(I)$ a boundary face (with respect to the fixed disk $D_{0}$). Now consider two elementary intervals $I,I^{\prime}$. We will show that if $\Lambda(I)=\Lambda(I^{\prime})$ and $f(I)$ and $f(I^{\prime})$ are both maximal, then $f(I)=f(I^{\prime})$. Thus the number of maximal boundary faces is bounded by the number of distinct component sets that can occur, showing that the number of maximal faces is bounded. We need the following lemma.
Lemma 3.1.
Let $I,I^{\prime}$ be two elementary intervals such that $\Lambda(I)=\Lambda(I^{\prime})$ and let $D_{i},D_{j}\neq D_{0}$ be distinct disks that both contain $I$ and $I^{\prime}$. Then there is a curve $\gamma_{ij}\subset D_{0}\cap D_{i}\cap D_{j}$ that connects $I$ to $I^{\prime}$.
Proof 3.2.
Let $p$ be a point in the interior of $I$ and let $p^{\prime}$ be a point in the interior of $I^{\prime}$. Because $\Lambda(I)=\Lambda(I^{\prime})$, the faces $f(I)$ and $f(I^{\prime})$ must be contained in the same component of $D_{0}\cap D_{i}$. Hence, there is a curve $\gamma_{i}\subset D_{0}\cap D_{i}$ connecting $p$ to $p^{\prime}$. Similarly, there is a curve $\gamma_{j}\subset D_{0}\cap D_{j}$ connecting $p$ to $p^{\prime}$. Finally, because $I$ and $I^{\prime}$ are contained in the same component of $D_{i}\cap D_{j}$ there is a curve $\gamma^{*}\subset D_{i}\cap D_{j}$ connecting $p$ to $p^{\prime}$. The curve $\gamma^{*}$ may lie partially outside $D_{0}$. Every time $\gamma^{*}$ exits $D_{0}$, it must re-enter $D_{0}$ at some point on $\partial D_{0}$. Thus we can replace the portions of $\gamma^{*}$ that lie outside $D_{0}$ by paths that follow $\partial D_{0}$; see Fig. 6(i). We denote the rerouted curve by $\gamma^{\prime}$. By construction, $\gamma^{\prime}$ connects $p$ to $p^{\prime}$ and is contained in $D_{0}$. We claim that $\gamma^{\prime}$ is also contained in $D_{i}$ and in $D_{j}$. Thus, we can take $\gamma_{ij}:=\gamma^{\prime}$ to prove the lemma.
Figure 6: (i) The curve $\gamma^{*}$ is rerouted to obtain a curve $\gamma^{\prime}$ (shown in orange) staying inside $D_{0}$. (ii) The left and right envelope (shown in orange) of a set of curves (shown in green).
Next we prove our claim that $\gamma^{\prime}$ is contained in $D_{i}$; the proof that it is contained in $D_{j}$ is analogous. Consider the closed curve $\gamma_{i}\cup\gamma^{*}$, which is contained in $D_{i}$. Since $D_{i}$ is a topological disk, the region $R^{*}$ enclosed by the (possibly self-intersecting) curve $\gamma_{i}\cup\gamma^{*}$ must be contained in $D_{i}$. Note that the region $R^{\prime}$ enclosed by $\gamma_{i}\cup\gamma^{\prime}$ is contained in $R^{*}$; indeed, the replacement of $\gamma$ by $\gamma^{\prime}$ removes the portions of $R^{*}$ that lie outside $D_{0}$. Thus, $R^{\prime}$ and, hence, $\gamma^{\prime}$, are contained in $D_{i}$, as claimed.
We can now prove that any two elementary intervals with the same component set that both bound a maximal boundary face, must in fact bound the same face.
Lemma 3.3.
Let $I,I^{\prime}\in\mathcal{I}$ be two elementary intervals with $\Lambda(I)=\Lambda(I^{\prime})$ and such that the boundary faces $f(I)$ and $f(I^{\prime})$ are both maximal. Then $f(I)=f(I^{\prime})$.
Proof 3.4.
If $I$ and $I^{\prime}$ are contained in a single disk $D_{i}\neq D_{0}$, then the lemma is easily seen to hold. Indeed, since $\Lambda(I)=\Lambda(I^{\prime})$, we know that $I$ and $I^{\prime}$ are contained in the same component of $D_{0}\cap D_{i}$, and so there is a curve $\gamma_{i}$ connecting $I$ to $I^{\prime}$ that stays inside this component. The curve $\gamma_{i}$ cannot intersect the boundary of any disk $D_{j}\not\in\{D_{0},D_{i}\}$, since the first time $\gamma_{i}$ would intersect such a boundary it would enter the corresponding disk $D_{j}$—this is true because $D_{j}$ does not contain II—which contradicts the maximality of $f(I)$. So from now on we assume that $I$ and $I^{\prime}$ are contained in at least two disks, in addition to $D_{0}$. Recall that each component $C\in\mathcal{C}(\mathcal{D})$ is a connected component of the intersection $D_{i}\cap D_{j}$ defined by some pair $D_{i},D_{j}\in\mathcal{D}$. Now consider a disk $D_{i}\neq D_{0}$, and define $\Lambda_{i}(I)\subset\Lambda(I)$ to be the set of components for which $D_{i}$ is one of the defining disks. In other words,
[…]
For each component $C\in\Lambda_{i}(I)$ of $D_{i}\cap D_{j}$ there is a curve $\gamma_{ij}\subset D_{0}\cap D_{i}\cap D_{j}$ that connects $I$ to $I^{\prime}$, by Lemma 3.1. Fix one such curve for each pair $(i,j)$, and let $\Gamma_{i}$ denote the collection of all these curves $\gamma_{ij}$. Note that any two sets $\Gamma_{i},\Gamma_{j}$ have a curve in common, namely $\gamma_{ij}$. The part of $\partial D_{0}$ not covered by $I$ or $I^{\prime}$ consists of two intervals. Let $p_{\mathrm{left}}$ be a point in the interior of one of these intervals and $p_{\mathrm{right}}$ to be a point in the interior of the other interval. We use the points $p_{\mathrm{left}}$ and $p_{\mathrm{right}}$ to define the left and right envelope of a given collection $\Gamma$ of curves connecting $I$ to $I^{\prime}$, as follows. Consider the arrangement $\mathcal{A}:=\mathcal{A}(\Gamma\cup\{\partial D_{0}\})$ defined by the curves in $\Gamma$ and $\partial D_{0}$ within $D_{0}$. Let $f_{\mathrm{left}}$ be the face of $\mathcal{A}$ containing $p_{\mathrm{left}}$, and let $f_{\mathrm{left}}$ be the face of $\mathcal{A}$ containing $p_{\mathrm{left}}$. Then the left envelope of $\Gamma$, denoted by $\mathcal{E}_{\mathrm{left}}(\Gamma)$, is the part of the boundary of $f_{\mathrm{left}}$ lying strictly inside $D_{0}$. Similarly, the right envelope $\mathcal{E}_{\mathrm{right}}(\Gamma)$ is is the part of the boundary of $f_{\mathrm{right}}$ lying strictly inside $D_{0}$; see Fig. 6(ii) for an illustration. Now consider the left and right envelopes $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ and $\mathcal{E}_{\mathrm{right}}(\Gamma_{i})$ of the sets $\Gamma_{i}$ defined above. Let $R_{i}$ be the region in $D_{0}$ between $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ and $\mathcal{E}_{\mathrm{right}}(\Gamma_{i})$. Since $D_{i}$ is a topological disk, and all curves in $\Gamma_{i}$ as well as $I$ and $I^{\prime}$ are contained in $D_{i}$, we know that $R_{i}\subset D_{i}$. Let
$B_{\mathrm{left}}:=\{\mathcal{E}_{\mathrm{left}}(\Gamma_{i}):I\subset D_{i}\}$
denote the set of left envelopes, and define
$Z_{\mathrm{left}}:=\mathcal{E}_{\mathrm{right}}(B_{\mathrm{left}})$
to be the right envelope of the left envelopes. Similarly, define $Z_{\mathrm{right}}$ to be the left envelope of the set $B_{\mathrm{right}}:=\{\mathcal{E}_{\mathrm{right}}(\Gamma_{i}):I\subset D_{i}\}$ of right envelopes. We will now consider the relative position of $Z_{\mathrm{left}}$ and $Z_{\mathrm{right}}$ to prove that there must be a curve $\gamma^{*}$ connecting $I$ to $I^{\prime}$ that lies in the common intersection of the disks $D_{i}$ that contain $I$. In other words, we can go from $I$ to $I^{\prime}$ via $\gamma^{*}$ without leaving any of the disks containing $I^{\prime}$. Now, if $f(I)$ is a maximal face, then $\gamma^{*}$ cannot enter any disk, thus proving that $f(I)=f(I^{\prime})$. Hence, to prove the lemma, it remains to argue the existence of $\gamma^{*}$. We consider three cases. (In these cases, the terms left and right are defined using the points $p_{\mathrm{left}}$ and $p_{\mathrm{right}}$, as before. For instance, $Z_{\mathrm{left}}$ lying completely to the left of $Z_{\mathrm{right}}$ means that $Z_{\mathrm{left}}$ lies in the region bounded by $Z_{\mathrm{right}}$ and the part of $\partial D_{0}$ that contains $p_{\mathrm{left}}$.)
  • Case I: $Z_{\mathrm{left}}$ lies completely to the left of $Z_{\mathrm{right}}$.
    Let $\gamma^{*}$ be any curve between $Z_{\mathrm{left}}$ and $Z_{\mathrm{right}}$, and consider any disk $D_{i}$ containing $I$. By definition of $Z_{\mathrm{left}}$, the curve $\gamma^{*}$ lies to the right of all left envelopes. In particular, $\gamma^{*}$ lies to the right of $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$. Similarly, $\gamma^{*}$ lies to the left of all right envelopes, and so $\gamma^{*}$ lies to the left of $\mathcal{E}_{\mathrm{right}}(\Gamma_{i})$. Hence, $\gamma^{*}\subset R_{i}\subset D_{i}$. Since this holds for any disk $D_{i}$ containing $I$, we conclude that $\gamma^{*}$ has the desired properties.
  • Case II: $Z_{\mathrm{left}}$ lies completely to the right of $Z_{\mathrm{right}}$.
    We claim that this cannot happen. To see this, let $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ define the endpoint of $Z_{\mathrm{left}}$ on $I$, and let and $\mathcal{E}_{\mathrm{right}}(\Gamma_{j})$ define the endpoint of $Z_{\mathrm{right}}$ on $I$. Then $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ lies to the right of $\mathcal{E}_{\mathrm{right}}(\Gamma_{j})$ along $I$. But this is not possible because there is a curve, $\gamma_{ij}$, that is a curve in $\Gamma_{i}$ and in $\Gamma_{j}$. Thus along $I$ the leftmost curve of $\Gamma_{i}$ cannot lie to the right of the rightmost curve in $\Gamma_{j}$.
  • Case III: $Z_{\mathrm{left}}$ crosses $Z_{\mathrm{right}}$.
    We will prove that this cannot happen either. To this end, let $x$ be an crossing point of $Z_{\mathrm{left}}$ and $Z_{\mathrm{right}}$. Let $Z_{\mathrm{left}}$ and $Z_{\mathrm{right}}$ be defined by $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ and $\mathcal{E}_{\mathrm{right}}(\Gamma_{j})$ at the point $x$, respectively, so that $x\in\mathcal{E}_{\mathrm{left}}(\Gamma_{i})\cap\mathcal{E}_{\mathrm{right}}(\Gamma_{j})$. Note that the endpoint of $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ on $I$ must lie to the left of the endpoint of $\mathcal{E}_{\mathrm{right}}(\Gamma_{j})$ on $I$; this follows by the same reasoning as in Case II. The curve $\gamma_{ij}\in\Gamma_{i}\cap\Gamma_{j}$ starts between $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ and $\mathcal{E}_{\mathrm{right}}(\Gamma_{j})$ on $I$, and by definition it reaches $I^{\prime}$ while staying inside $R_{i}\cap R_{j}$. But this is impossible if $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ and $\mathcal{E}_{\mathrm{right}}(\Gamma_{j})$ cross, because then $\gamma_{ij}$ must cross at least one of these curves, say $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$, in order to reach $I^{\prime}$. This is a contradiction, because if $\gamma_{ij}$ crosses $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$ for the first time, it would enter the part of $D_{0}$ to the left of $\mathcal{E}_{\mathrm{left}}(\Gamma_{i})$, which contradicts the definition of $\mathcal{E}_{\mathrm{left}}(\gamma_{i})$ since $\gamma_{ij}\in\Gamma_{i}$.
We are now in a position to prove Theorem 1.2, which states that $\mu(\mathcal{A})$, the number of maximal faces in the arrangement induced by a set $\mathcal{D}$ of $n$ topological disks with overlap number $\Delta$, is bounded by $n(\Delta+1)^{n(n-1)/2}$. To this end, observe that the number of maximal boundary faces of a fixed disk $D_{i}\in\mathcal{D}$ is bounded by the number of distinct component sets $\Lambda(I)$, by Lemma 3.3. Also observe that a component set $\Lambda(I)$ contains at most one component from each set $\mathcal{C}(D_{j},D_{k})$, and that $|\mathcal{C}(D_{j},D_{k})|\leqslant\Delta$ by the definition of $\Delta$. Hence, the number of different component sets is bounded by $(\Delta+1)^{n(n-1)/2}$. Multiplying this by $n$, the number of different choices for the fixed disk $D_{i}$, we obtain the desired bound of $n(\Delta+1)^{n(n-1)/2}$ on the number of maximal faces.
3.2 The proof of Theorem 1.3
@@
Using the bound on the number of maximal faces in the arrangement $\mathcal{A}$, we now obtain a bound on the diameter of the dual graph $\mathcal{G}^{*}$ of $\mathcal{A}$. Recall that the ply of a face $f$ in the arrangement $\mathcal{A}$, denoted by $\operatorname{ply}(f)$, is the total number of disks in $\mathcal{D}$ that contain $f$. Define $p_{\max}:=\max_{f}\operatorname{ply}(f)$ to be the maximum ply over all faces in the arrangement $\mathcal{A}$. For a node $u$ in the dual graph $\mathcal{G}^{*}$, we define $\operatorname{ply}(u):=\operatorname{ply}(f_{u})$, where $f_{u}$ is the face in $\mathcal{A}$ corresponding to $u$. The next lemma relates $\mu(\mathcal{A})$, the number of maximal faces in $\mathcal{A}$, to $\operatorname{diam}(\mathcal{G}^{*})$, the diameter of the dual graph $\mathcal{G}^{*}$ of $\mathcal{A}$. Combining the lemma with Theorem 1.2 proves Theorem 1.3.
@@
Note that each node $u\in V$ is assigned to a set $V_{i}$, because from $u$ we can greedily move to faces of higher ply until we reach a maximal face. Also note that $v_{i}\in V_{i}$ for all $i$. Now, let $\overline{\mathcal{G}^{*}}$ be the graph obtained from $\mathcal{G}^{*}$ by contracting each subgraph $\mathcal{G}^{*}[V_{i}]$ to a single node. We make the following two observations.
@@
Figure 7: An example with $\Theta(n^{2}\Delta)$ maximal faces: in $n\times n$ grid for every grid cell on the main diagonal we introduce a disk inscribed in the interior of the cell having $k$ parallel strains extended to the four sides of the grid. By construction, there are $n$ disks and $\Delta=2k^{2}$ while the number of maximal faces is $n(n-1)k^{2}=n(n-1)\Delta/2$.
@@ 4 Concluding Remarks and Discussion
In this paper, we derive bounds on the diameter of the dual graph of an arrangement $\mathcal{A}$ of disks in terms of the number of disks, $n$, and the maximum number of intersection components between any two disks, $\Delta$. For an arrangement containing only two disks, we establish a sharp linear bound of $\max\{2,2\Delta\}$. Such upper bound is equivalent to a bound on the number of crossings of any minimum-crossing $st$-curve in arrangements of two disks. For an arrangement $\mathcal{A}$ of $n$ disks, we bound both the diameter of the dual graph and the number of maximal faces as a function of $n$ and $\Delta$. The bound on the number of maximal faces $\mu(\mathcal{A})$ is $n(\Delta+1)^{n(n-1)/2}$. Therefore the upper bound on the diameter of the dual graph of arrangements with $n$ disks is $2n(\Delta+1)^{n(n-1)/2}\min\{n,\Delta+1\}$, which corresponds to the bound for the number of crossings of any minimum-crossing $st$-curve. We end this paper with an open problem: Is it possible to derive an upper bound on $\mu(\mathcal{A})$ which is polynomial in $n$ and $\Delta$? Our best example has $\Theta(n^{2}\Delta)$ maximal faces, see Fig. 7. We conjecture that this bound for $\mu(\mathcal{A})$ is tight up to a constant factor.

### v2

@@ 3 The general case of $n$ disks
Using the bound on the number of maximal faces in the arrangement $\mathcal{A}$, we now obtain a bound on the diameter of the dual graph $\mathcal{G}^{*}$ of $\mathcal{A}$.
@@
Recall that the ply of a face $f$ in the arrangement $\mathcal{A}$, denoted by $\operatorname{ply}(f)$, is the total number of disks in $\mathcal{D}$ that contain $f$. Define $p_{\max}:=\max_{f}\operatorname{ply}(f)$ to be the maximum ply over all faces in the arrangement $\mathcal{A}$. For a node $u$ in the dual graph $\mathcal{G}^{*}$, we define $\operatorname{ply}(u):=\operatorname{ply}(f_{u})$, where $f_{u}$ is the face in $\mathcal{A}$ corresponding to $u$. The next lemma relates $\mu(\mathcal{A})$, the number of maximal faces in $\mathcal{A}$, to $\operatorname{diam}(\mathcal{G}^{*})$, the diameter of the dual graph $\mathcal{G}^{*}$ of $\mathcal{A}$. Combining the lemma with Theorem 1.3 proves Theorem 1.4.
@@
Note that each node $u\in V$ is assigned to a set $V_{i}$, because from $u$ we can greedily move to faces of higher ply until we reach a maximal face. Also note that $v_{i}\in V_{i}$ for all $i$.
Now, let $\overline{\mathcal{G}^{*}}$ be the graph obtained from $\mathcal{G}^{*}$ by contracting each subgraph $\mathcal{G}^{*}[V_{i}]$ to a single node. We make the following two observations.
@@ 4 Concluding Remarks and Discussion
In this paper, we derived bounds on the diameter of the dual graph of an arrangement $\mathcal{A}$ of disks in terms of the number of disks, $n$, and the maximum number of intersection components between any two disks, $\Delta$. (The bound on the diameter is equivalent to a bound on the number of crossings of any minimum-crossing $st$-curve in arrangements of two disks.)
For an arrangement containing only two disks, we established a sharp linear bound of $\max\{2,2\Delta\}$. For an arrangement of $n$ disks, we bounded the diameter of the dual graph by $O(n^{3}2^{n}\Delta)$. This was done by establishing an $O(n^{2}2^{n}\Delta)$ bound on the number of maximal faces in $\mathcal{A}$. The latter was done by first proving an $O(n^{2}\Delta)$ bound on the number of maximum faces, which is asymptotically tight.
Our bound on the number of maximal faces is far from tight: our best lower bound is the same as for the number of maximum faces, namely $\Omega(n^{2}\Delta)$, while our upper bound is $O(n^{2}2^{n}\Delta)$. The main open problem is to tighten this gap. We believe that the true bound is polynomial in nn—in fact, we conjecture that the $\Omega(n^{2}\Delta)$ lower bound is asymptotically tight. The upper bound on the number of maximal faces was used to derive a bound on the diameter of the dual graph of the arrangement. Notice that by making $n/2$ nested copies of each of the two disks in Figure 4, we can produce an arrangement of $n$ disks with overlap $\Delta$ and diameter $\Omega(n\Delta)$. Instead of trying to improve the $O(n^{3}2^{n}\Delta)$ upper bound by improving the bound on the number of maximal faces, one may also try to prove a better bound directly.